OpenStudy (anonymous):

HELP PLEASE!!!! locate the critical points and classify them as stationary points. f(x)=sin^2x, 0

OpenStudy (anonymous):

Do you know how to find critical points?

OpenStudy (anonymous):

To find critical points you take the derivative and set it to 0 and solve for x, so take the derivative of your problem then set it = 0 and then solve for x.

OpenStudy (rajee_sam):

\[Sin ^{2}X\]

OpenStudy (rajee_sam):

is it ?

OpenStudy (anonymous):

oh no its sin^2 2x

OpenStudy (rajee_sam):

OK to find the critical points like onegirl said set your first derivative to zero and find the values of x for which it will be = 0

OpenStudy (rajee_sam):

\[f'(x) = 2\sin2xcos2x . 2 = 4 \sin2x \cos2x\]

OpenStudy (rajee_sam):

\[4 \sin2xcos2x = 0\] means either sin2x = 0 or cos2x = 0

OpenStudy (rajee_sam):

your interval given is 0 to 2pi

OpenStudy (anonymous):

how did you get that first derivative?

OpenStudy (anonymous):

was it with a chain rule?

OpenStudy (rajee_sam):

yes

OpenStudy (rajee_sam):

sin^2 2x - first do the power 2 sin2x, now do the sin2x, which will cos2x and then you have to do the 2x which is 2

OpenStudy (anonymous):

oh okay i see now

OpenStudy (rajee_sam):

 If you see the curve slope is zero at \[\pi/4 , 3\pi/4, 5\pi/4, 7\pi/4\]